OpenStudy (anonymous):

Isolate y in the equation 3x + 2y = 8. A y=-3/2x-4 B y=-3/2x+4 c y=3/2x+4 D y=3/2x-4

OpenStudy (anonymous):

2y = 8 - 3x y = 4 - 3/2 x